Please help. The line passing though (7,y) and (1,0) is parallel to the line joining (4,4)and (-2,1). Find y.
Found 2 solutions by rfer, lwsshak3:
Answer by rfer(16322)

m=(1-4)/(-2-4)
m=-3/-6
m=1/2
------------------
y=mx+b
0=1/2*1+b
b=-1/2
---------------
y=1/2*7+1/2
y=7/2+1/2
y=8/2
y=4

Answer by lwsshak3(11628)

Lines parallel to each other have the same slope
slope of line joining (4,4) and (-2,1) =∆y/∆x =(4-1)/(4-(-2))=3/6=1/2
for line passing thru (7,y) and (1,0), using point (1,0) and slope intercept formula, y = mx + b
0=(1/2)*1 + b
b=-(1/2)
equation of line passing thru (7,y) and (1,0)
y=(1/2)x-(1/2)
when x=7, y = (1/2)*7-(1/2) = 7/2 -1/2 =6/2 =3
Ans: y=3